# Aligning Agents like Large Language Models

### v1

@@ 3.3 Obtaining Preference Data on Online Rollouts
We now deploy the fine-tuned agent in the environment to collect trajectories for preference labelling. This is analogous to generating multiple LLM responses to a prompt, but in this context the prompt becomes the initial observation and optional context of any previous observations and actions provided. Similarly to LLMs, there are tradeoffs to be made in the diversity, quality and quantity of preferences when provided by humans (discussed further in Appendix I), but here we avoid such issues by utilizing synthetic preferences to isolate the effect of quantity.
@@
Online Rollouts: We initialize the Ninja character with an empty context buffer at a random spawn point, equivalent to the evaluation procedure above. We repeated this procedure to generate a video dataset of 2400 on-policy trajectories, divided into 1000 trajectories for training and 1400 for evaluation. Similarly to LLMs, the temperature of the softmax sampling of the policy for action selection can be increased to generate more diverse behaviors from the agent for easier comparison, but here we found the behavior to be diverse enough with a default temperature parameter of 1.
@@
Generating Preferences: We utilize synthetic preferences based on the primary criteria:
Preferred Jumppad Reached $>$ Other Jumppad Reached $>$ No Jumppad Reached
@@
Within each of these primary categories, we further rank trajectories by their duration, with shorter trajectories being preferred. By selecting subsets of trajectories in the training dataset, we are able to investigate how reward model performance scales with number of comparisons (which is a proxy for the human labelling time requirement, often the main bottleneck for reinforcement learning from human feedback, or RLHF). A preference dataset $P$ is then created by considering all pairwise comparisons of trajectories $\tau$ using the criteria above, i.e. $(\tau_{A}\succ\tau_{B})\in P\ \forall\ \tau_{A},\tau_{B}\in\tau\text{ if }\tau_{A}\succ\tau_{B}$.
3.4 Training a Reward Model on Preferences
@@
A reward model is then trained on this dataset to predict rewards in accordance with preferences as described below. Previous work on RLHF for agents has generally relied on simple (often linear) reward models (Knox & Stone, 2008; Christiano et al., 2017). However recent work on LLMs has demonstrated that reward models that utilize the pre-trained or fine-tuned policy model with the action classification head replaced with a scalar regression head generally perform better (Stiennon et al., 2020), and also improve with scale (Ouyang et al., 2022; Touvron et al., 2023). It is hypothesised that this is because this allows the reward model to benefit from the pre-training, which helps to the reward model to better distinguish behaviors for reward assignment, since the reward model shares the agent’s knowledge of the behavior space (Ziegler et al., 2020). We investigate this potential for representation transfer in the context of agents in this section.
@@
Procedure: We follow the standard Bradley-Terry (Bradley & Terry, 1952) model procedure to train a reward model $\hat{r}$ from these pairwise preferences, as introduced by Christiano et al. (2017). Specifically, we interpret trajectory rewards as preference rankings analogous to Elo ratings (Elo, 1978) developed for chess, such that the annotator’s probability of preferring a trajectory depends exponentially on the trajectory reward difference. We can then fit the reward model by minimising the cross-entropy between these probabilities and the preference labels, which gives the loss function:
@@
where $\sigma$ is the sigmoid function and $(\tau_{w},\tau_{l})$ are the trajectories being compared, with $\tau_{w}$ being the winning (preferred) trajectory and $\tau_{l}$ being the losing trajectory. Since the reward model is only trained on comparisons, the scale of the predicted rewards is arbitrary. As a result, we found the need to apply a small amount of $L2$ regularisation to prevent the scale of the rewards becoming overly large for the best and worst trajectories (overfitting). We further empirically normalise the reward model after training using the max and min of predicted rewards over the training trajectories to scale our reward model output to be in the range $\hat{r}\in[0,1]$ (a crucial but often overlooked detail).
@@
Architecture: Following the LLM procedure to use the agent model for the reward model initialization, we take the outputs of the fine-tuned agent’s transformer as embeddings for each timestep, which are fed into an MLP reward model to produce a scalar return for the trajectory which is then input to the loss function in Equation 1. As an ablation, to investigate whether these observation representations used by the agent for imitation learning are also beneficial for reward modelling (i.e. predicting preferences), we consider a reward model with an equivalent architecture but with a randomly initialized linear encoder to provide random projections as embeddings for each timestep. We also investigate how reward model performance scales with number of comparisons, to obtain an estimate of the human time required to provide sufficient feedback for training the reward model.
Figure 5: Reward model test performances.
@@
Results: To measure the performance of our reward models, we apply each reward model to our held-out test trajectories. We compute pairwise preferences according to the reward models, and compare to the ground truth pairwise preferences to obtain a test preference accuracy. Reward model performances shown in Figure 5.
@@
We find that reward model accuracy generally increases with number of comparisons, although the random projection reward models have high variance. Most strikingly, the reward models utilizing the agent model perform better than the reward models using random projections across the full range of comparison sizes, suggesting that the imitation learning agent’s representations of the observations contains information beneficial to predicting preferences. We further find that when using this encoder it is possible to train a reward model from visual input to achieve over 90% preference accuracy with only $\sim 100$ comparisons. While we use synthetic preferences, we note that this would correspond to less than 1 hour of labelling time for our task.
3.5 Aligning the Fine-tuned Model with the Reward Model
Procedure: Finally, we align our agent with our preferences as captured by our reward models. We run our fine-tuned agent in the environment as described in Section 2 to generate online trajectories. After each trajectory is complete, we apply our reward model to the trajectory to generate a reward corresponding to that trajectory. We use this reward as the return for that trajectory and update our agent policy $\pi$ using an undiscounted REINFORCE (Williams, 1992) loss. While much of the RLHF literature uses PPO (Schulman et al., 2017), we note that concurrent work on LLMs has found PPO to be unnecessarily complex for RLHF (Ahmadian et al., 2024). This is due to the fact that in the standard LLM framework the reward is only provided at the end of the trajectory, meaning that we generally just want to reinforce entire trajectories rather than attempt credit assignment at each timestep using an advantage function. Additionally, Ahmadian et al. (2024) found that clipping rarely occurs in practice since the initial fine-tuned policy is close to the final aligned policy. This is reflected by the use of REINFORCE in recent state-of-the-art LLMs (Google Deepmind, 2024).
@@ 3.5 Aligning the Fine-tuned Model with the Reward Model
For our experiments we ran our agent online for 9600 episodes (corresponding to around 1 day of real time gameplay), using a batch size of 16 to give 600 parameter updates. We fine-tune the last layer only since we want to maintain the general pre-trained representations, but note that approaches such as Low-Rank Adaption (LoRA) (Hu et al., 2021) used for fine-tuning LLMs could also be used here. We also investigated the use of an additional KL divergence term to regularise the optimized policy towards the initial fine-tuned policy, as is commonly used in LLM alignment (Touvron et al., 2023; Bai et al., 2022), but found it unnecessary. This is likely due to the fact that we fine-tune only the last layer rather than the entire network, which mitigates policy divergence.
@@ 3.5.1 Aligning Agent Towards Left Jumppad
Figure 6: Left jumppad success rate during online alignment. Standard errors shaded, and number of seeds shown in legend. We see that: 1) Higher accuracy reward models (darker) generally lead to better alignment, 2) Preference fine-tuning (blue) on preferred trajectories before online alignment improves performance for all reward models.
We begin by focusing on aligning our agent towards the left jumppad. We plot the average success rate of reaching the left jumppad against the number of episodes with reward models that have been trained on 100 up to 500k comparisons in Figure 6. We see that all of our reward models are sufficient to align our agent to consistently reach the left jumppad, with reward models trained on more data (with higher test performances) generally leading to better alignment.
However, we see that the agents take most of the 600 updates to fully align, corresponding over a day of real time training. To improve this efficiency, we consider the addition of a preference fine-tuning phase in which we first further fine-tune on the preferred trajectories. Specifically, we apply the reward model to the training trajectories and take the top 20% of trajectories by reward (corresponding to the preferred trajectories) and perform additional fine-tuning on these trajectories. This is similar to Reinforced Self-Training (ReST) (Gulcehre et al., 2023), recently introduced to improve the efficiency of language model alignment. We find that this improves online performance for all reward models, shown in Figure 6.
@@ 3.5.2 Aligning Agent Towards Right Jumppad
Figure 7: Right jumppad success rate during online alignment. Standard errors shaded, and number of seeds in legend. Similar trends hold to Figure 6, but alignment is less effective.
In order to investigate this discrepancy, we analyzed the rollouts of the fine-tuned agent, shown in Figure 8. We see that the trajectories that reach the left jumppad start relatively evenly from the 4 spawn locations. However, of the trajectories that successfully reach the right jumppad, only 1% originated from the spawns on the right hand side. Therefore it is unsurprising that it is more difficult for the agent to learn to go right, given that in half the episodes (from the right spawn locations) the agent rarely navigates to the right jumppad to receive reward to reinforce its behavior. To confirm this hypothesis, we instead attempted to align the base agent (before fine-tuning), and found that it can indeed be aligned more symmetrically, although with a lower final success rate, as discussed in Appendix 17. The addition of preference fine-tuning before online alignment helps to mitigate this imbalance, as shown in Figure 8, since it modifies the agent to produce more of the preferred behaviors that achieve reward before online alignment. This enabled us to completely align the agent right with the same training budget, as shown by the final jumppad distributions in Appendix 19. This further demonstrates the benefits of our proposed additional stage of preference fine-tuning.
@@ 3.6 Summary of Agent Alignment
Figure 9: Heatmap of agent trajectories at each stage of our pipeline. Each stage shows 1000 rollouts.
To visualise the gradual alignment of our agent, we provide a heatmap of the agent trajectories at each stage of our alignment pipeline above in Figure 9. We see that the base agent trajectories are relatively diverse, capturing a variety of human behaviors, including those that do not reach a jumppad, or take an indirect route to do so. The fine-tuned agent which has been refined on curated task-specific/demonstration trajectories shows more direct trajectories to the jumppads, but still does not incorporate any preference regarding the the jumppad which we would like our agent to reach. Finally, preference fine-tuning and subsequent online reinforcement learning enable us to align our agent to reliably perform the desired behavior; in this case to reach the left or the right jumppad. Videos of the behavior of our agent at each stage of this alignment pipeline are provided on the project webpage at: https://adamjelley.github.io/aligning-agents-like-llms/.
4 Related Work
Large Scale Imitation Learning: Imitation learning and offline reinforcement learning (Levine et al., 2020) have recently gained popularity with the aim of demonstrating similar scaling laws to LLMs (Kaplan et al., 2020; Hernandez et al., 2021) to obtain more general agents. Decision Transformer (Chen et al., 2021a) proposed imitation learning with a transformer policy that can be conditioned on a desired return. Multi-Game Decison Transformers (Lee et al., 2022) extended this approach to learn multiple game policies with a single model. GATO (Reed et al., 2022) and RoboCat (Bousmalis et al., 2023) demonstrated large scale multi-task imitation learning could enable a single model to learn and improve on hundreds of diverse tasks. More recently, SIMA (SIMA-Team et al., 2024) is a general agent with language input and keyboard output which shows promising generalization to new video games, and further demonstrates the potential for scaling imitation learning to large, diverse datasets.
Fine-tuning with Reinforcement Learning: The use of imitation learning as pre-training for reinforcement learning was first investigated to improve the sample efficiency of deep RL algorithms by reducing the exploration space (Hester et al., 2017; Vecerik et al., 2018). AlphaGo (Silver et al., 2016) and subsequently AlphaStar (Vinyals et al., 2019) demonstrated that imitation learning on human data could provide a strong behavior prior for environments like StarCraft where reinforcement learning from scratch is infeasible. VPT (Baker et al., 2022) extended this paradigm to web-scale data by performing imitation learning on 70k hours of Minecraft videos labelled by an inverse dynamics model, before fine-tuning with reinforcement learning on task-specific rewards. While these works demonstrate the potential of fine-tuning large imitation models, they use hard-coded reward functions to maximise agent performance rather than align an agent’s behavior with subjective preferences.
Reinforcement Learning from Human Feedback (RLHF): Training agents with human preferences has a long history, as reviewed by Wirth et al. (2017) and Zhang et al. (2021) (including notably Bennett et al. (2009) and Knox & Stone (2008)), leading to the popular modern formulation of RLHF for deep learning proposed by Christiano et al. (2017) that we use in our work (see Section 3.4). This idea was extended by Ibarz et al. (2018) to include imitation learning as pre-training to improve the efficiency of preference learning. PEBBLE (Lee et al., 2021) instead utilises unsupervised pre-training to increase the diversity of initial behaviors. Abramson et al. (2022) and Milani et al. (2023) are closest to our work in demonstrating the potential of RLHF for improving the performance of large imitation agents in 3D simulated worlds, but do not analyze the full LLM training pipeline.
Large Language Models: Radford et al. (2018) demonstrated that pre-training with an unsupervised generative task (next token prediction) on a large diverse corpus of text, followed by fine-tuning on a specific task provided benefits compared fine-tuning alone. Stiennon et al. (2020) (following Ziegler et al. (2020)) then popularised the use of RLHF to further fine-tune these models to align their responses with human preferences. This procedure led to InstructGPT (Ouyang et al., 2022), Chat-GPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023) as well as open-source models, such as Claude (Bai et al., 2022), Llama 2 and 3 (Touvron et al., 2023) and Gemma (Google Deepmind, 2024). The success of these models suggests that it is worth investigating whether their training procedure can be transferred to other domains. While components of our procedure have been applied both individually and in combinations to agents, our work is the first to investigate the benefits of applying the full LLM training procedure to agents trained end to end from visual inputs to a unified action space.
5 Limitations and Broader Impact
One limitation of our work is that we assume access to a large amount of human data on the game of interest to pre-train a base model. However, large amounts of human data may not be available. A potential solution may be to use a ‘gaming foundation model’ trained on similar games (with pixel input and action controller output), as investigated by other works (Lee et al., 2022; Reed et al., 2022; SIMA-Team et al., 2024). Another limitation is our use of synthetic preferences. While this may be possible for some games, real human feedback from a game designer will be required in general, and may introduce label noise. A final limitation is that running agents online in modern games to collect preferences and for fine-tuning can be expensive, especially since these games must often be run in real-time in order to be rendered. Preference fine-tuning, as we introduced, provides one way to improve the efficiency of alignment. Other developments in the use of RLHF for language models could also be applied to address these challenges, as discussed in Appendix I, and we believe could provide many promising research directions to improve the efficiency of agent alignment.
More generally, research into aligning agents with preferences is important to ensure that agents are helpful and harmless. However, this procedure has various known open problems and limitations (Casper et al., 2023). Video games therefore provide an important test bed for such research, helping to mitigate risks and provide insights in a safe environment that may generalize to other more real-world applications such as robotics.
6 Conclusion
In this paper, we have drawn an analogy between training agents and training large language models. We demonstrated that the current LLM training procedure can be used to align agents to reliably perform desired behaviors in complex environments on a modern console game. This enabled us to train an agent to achieve specific modes of behavior in the game that would be difficult to achieve reliably with imitation learning, reinforcement learning or RLHF alone. Our analysis shows that many of the recent developments in the current procedure for training LLMs, such as general pre-training and initializing reward models from the pre-trained agent, can be applied and have similar benefits for training agents. We also demonstrated improved alignment efficiency through the inclusion of an additional preference fine-tuning stage. We hope that our work encourages further communication and collaboration between the LLM and gaming communities, to enable shared insights and provide a path towards practical and reliable deployment of reinforcement learning agents in modern games.
Acknowledgments and Disclosure of Funding

### v2

@@ 3.4 Preference Modeling
The next stage in the standard procedure for training LLMs is to obtain human preference data on model responses. Here we deploy the fine-tuned agent in the environment to collect trajectories for preference labelling. This is analogous to generating multiple LLM responses to a prompt, but in this context the prompt becomes the initial observation and optional context of previous observations and actions (or instructional text embeddings) provided. Similarly to LLMs, there are tradeoffs to be made in the diversity, quality and quantity of preferences (discussed further in Appendix J), but here we avoid such issues by utilizing synthetic preferences to isolate the effect of quantity. Additionally, this labeling could potentially be automated using recent VLMs, again leveraging ideas and developments from the LLM-as-a-Judge literature (zheng_judging_2023; li_llms-as-judges_2024).
@@
Online Rollouts: We use the evaluation procedure described in Section 3.2 to generate a dataset of 2400 on-policy trajectories, divided into 1000 trajectories for training and 1400 for evaluation. Similarly to LLMs, the temperature of the softmax sampling of the policy for action selection can be increased to generate more diverse behaviors from the agent for easier comparison, but here we found the behavior to be diverse enough with a default temperature of one.
@@
Generating Preferences: We use synthetic preferences to rank trajectories based on the primary jumppad (JP) criteria:
Given JP Reached $\succ$ Other JP Reached $\succ$ No JP Reached
@@
Within each of these categories, we further rank by duration, preferring shorter trajectories. By selecting subsets of the training trajectories, we can investigate how reward model performance scales with number of comparisons (which is a proxy for human labeling time, often the main bottleneck for RLHF). A preference dataset $P$ is then created by considering all pairwise comparisons of trajectories $\tau$ using the criteria above, i.e. $(\tau_{A}\succ\tau_{B})\in P\ \forall\ \tau_{A},\tau_{B}\in\tau\text{ if }\tau_{A}\succ\tau_{B}$.
3.5 Do modern reward-modeling practices apply in the context of agents?
@@
A reward model is then trained on this dataset to predict rewards in accordance with preferences, as described below. Previous work on RLHF for agents has generally relied on independent (often linear) reward models (knox_tamer_2008; christiano_deep_2017). However, recent work on LLMs has demonstrated that reward models using the pre-trained or fine-tuned policy model with the action classification head replaced with a scalar regression head generally perform better (stiennon_learning_2020), and also improve with scale (ouyang_training_2022; touvron_llama_2023). It is hypothesized that this is because the pre-training enables the reward model to better distinguish behaviors for reward assignment without the trajectory over-fitting that would occur with the equivalent model otherwise (ziegler_fine-tuning_2020). We investigate this potential for representation transfer in the context of visual input agents in this section.
@@
Procedure: We use the standard Bradley-Terry (bradley_rank_1952) modeling procedure to train a reward model $\hat{r}$ from these pairwise preferences, following christiano_deep_2017. Specifically, we interpret trajectory rewards as preference rankings analogous to Elo ratings (elo_rating_1978) developed for chess, such that the annotator’s probability of preferring a trajectory depends exponentially on the trajectory reward difference. We can then fit the reward model by minimising the cross-entropy between these probabilities and the preference labels, which gives the loss function:
@@
where $\sigma$ is the sigmoid function and $(\tau_{w},\tau_{l})$ are the trajectories being compared, with $\tau_{w}$ being the winning (preferred) trajectory and $\tau_{l}$ being the losing trajectory. Since the reward model is only trained on comparisons, the scale of the predicted rewards is arbitrary. As a result, we found the need to apply a small amount of $L2$ regularization to prevent the scale of the rewards becoming overly large for the best and worst trajectories. We further empirically normalize the reward model after training using the max and min of predicted rewards over the training trajectories to scale our reward model output to be in the range $\hat{r}\in[0,1]$.
@@
Architecture: Following the LLM procedure to use the agent model for the reward model initialization, we take the output embeddings of the fine-tuned transformer at each timestep as input to an MLP. This outputs a scalar return for use in the loss function in Equation 1. As a baseline, we considered training a reward model with an equivalent architecture end-to-end, but found this led to trivial over-fitting and negligible validation performance. Instead, we use a frozen randomly initialized encoder to provide embeddings for each timestep, which are fed into an equivalent MLP.
@@
Results: Each reward model is applied to our held-out test trajectories. We compute pairwise preferences, and compare to ground-truth preferences to obtain a test accuracy. Reward model performances are shown in Figure 7.
Figure 7: Reward model test accuracies against number of training comparisons for random and agent-initialized models.
@@
We find that reward model accuracy increases approximately logarithmically with comparisons, although the random projection reward models have high variance. Additionally, the reward models utilizing the agent model perform better than the reward models using random projections across the full range of comparison sizes, confirming the relevance of the imitation learning representations. Importantly, this transfer makes it possible to train a reward model from visual input to achieve over 90% preference accuracy with only $\sim 100$ comparisons. While this is a relatively simple task, we note that this would correspond to less than 1 hour of labeling time, demonstrating the feasibility of this approach for aligning pixel-based agents from trajectory videos.
In summary, we find evidence of representation transfer between imitation learning and preference modeling, that enables feedback-efficient reward modeling for aligning pixel-based agents.
3.6 Aligning the Agent with the Reward Model
Procedure: Finally, we align our agent with the preferences captured by our reward models. We run our fine-tuned agent in the environment as for evaluation to generate online trajectories. After each trajectory is complete, we apply our reward model to the trajectory to generate a reward. We use this reward as the return for the trajectory and update our agent policy $\pi$ using REINFORCE (williams_simple_1992). We discuss the motivation behind this choice in Appendix G.
@@ 3.6 Aligning the Agent with the Reward Model
For our experiments we ran our agent online for around a day of real-time gameplay, corresponding to 600 parameter updates (see Appendix C for details). The standard approach in LLM alignment would be to fine-tuning the entire network with an additional KL divergence term to regularize the optimized policy towards the initial fine-tuned policy (touvron_llama_2023; bai_training_2022), but for our proof of concept we simply fine-tune the last layer for efficiency. We also note that approaches such as Low-Rank Adaption (LoRA) (hu_lora_2021) used for fine-tuning LLMs could also be used here for memory-efficient alignment of a large base model with an individual’s preferences.
@@ 3.6.1 Aligning Agent Towards Left Jumppad
We begin by aligning our agent to reach the left jumppad. We plot the average success rate of reaching the left jumppad against the number of episodes in Figure 8, for reward models that have been trained on 100 up to 500k comparisons. We see that all of our reward models are sufficient to align our agent to consistently reach the left jumppad, with reward models trained on more data (with higher test performances) generally leading to better alignment.
Figure 8: Left jumppad success rate during online alignment. Standard errors shaded, and number of seeds shown in legend. We see: 1) Higher accuracy reward models (darker) generally lead to better alignment, 2) Preference fine-tuning before online alignment (blue) improves performance for all reward models.
However, we see that the agents take most of the 600 updates to fully align, corresponding to a day of real-time training. To improve this efficiency, we again take inspiration from the LLM literature and consider the addition of a preference fine-tuning phase in which we first further fine-tune on the preferred trajectories. Specifically, we apply the reward model to the training trajectories and take the top 20% of trajectories by reward (corresponding to the preferred trajectories) and perform additional fine-tuning on these trajectories. This corresponds to a single iteration of Reinforced Self-Training (ReST) (gulcehre_reinforced_2023), recently introduced to improve the efficiency of language model alignment. We find that this improves online performance for all reward models, shown in Figure 8.
@@ 3.6.2 Aligning Agent Towards Right Jumppad
Figure 9: Right jumppad success rate during online alignment. Standard errors shaded, and number of seeds shown in legend. Similar trends hold to Figure 8, with the benefits of preference fine-tuning (blue) even more prominent, but overall alignment is less effective.
We investigate this discrepancy in Appendix H, and find that it arises from an imbalance in spawn locations that makes it difficult for the agent to receive positive reward from certain spawns. However, the addition of preference fine-tuning before online alignment helps to mitigate this imbalance, as shown in Figure 9, since it modifies the agent to produce more of the preferred behaviors that achieve reward before online alignment. This enabled us to completely align the agent right with the same training budget, as shown by the final jumppad distributions in Appendix 20. This further demonstrates the benefits of our proposed additional stage of preference fine-tuning. However, this is just one possible strategy to improve alignment efficiency from the LLM literature. Many other approaches, such as Direct Preference Optimization (DPO) (rafailov_direct_2023) could also be incorporated to alleviate the need for costly on-policy, online alignment. However, removing the need for online learning entirely is likely to be challenging, as discussed in Appendix D, and has been recently demonstrated in the context of LLMs (xu_is_2024).
In summary, we find that agents can be aligned to behave in accordance with human preferences by leveraging equivalent approaches to those used for modern LLMs, with similar benefits and effectiveness.
@@ 3.7 Summary of Agent Alignment
To visualize the gradual alignment of our agent, we provide a heatmap of agent trajectories at each stage of the alignment pipeline in Figure 10. This demonstrates how pre-training, fine-tuning and RL(HF) can be used to reliably achieve behaviors that would not be possible with imitation learning (due to finite demonstrations) or reinforcement learning (due to lack of dense reward) alone. Videos of our agent at each stage are provided at: https://adamjelley.github.io/aligning-agents-like-llms.
@@
Figure 10: Heatmap of our agent’s trajectories at each stage of the alignment pipeline. Each stage shows 1000 trajectories.
4 Alternative Views and Broader Outlook
Our proof-of-concept provides evidence of the benefits of multi-stage training over imitation or reinforcement learning alone. However, an alternative view to our position might be that rather than training agents like LLMs, LLMs should be used as agents directly. There has indeed been significant progress in the development of LLM agents in recent years (xi_rise_2023; wang_survey_2024), including for low-level control (yang_towards_2024; brohan_rt-2_2023; ahn_as_2022; driess_palm-e_2023; tang_saytap_2023).
However, multi-modal LLMs still struggle with low-level decision-making in complex environments such as video games, with recent work demonstrating that these models are a long way from human performance in these environments, even with significant amounts of gameplay data provided for in-context learning (waytowich_atari-gpt_2024). In particular, paglieri_balrog_2024 recently demonstrated that LLM agents understandably struggle with experience-based aspects such as spatial reasoning, systematic exploration, long-term planning, understanding environment dynamics, and the ‘knowing-doing’ gap. In fact, LMAct (ruoss_lmact_2024) concludes “Our work thus adds another piece to the growing literature that finds difficulties that LMs can have with translating declarative knowledge into effective acting (know-how) in some interactive decision-making tasks.
We argue that for tasks such as playing a 3D AAA video game, requiring fine-grained control of a gamepad controller at 10Hz, this problem will not simply go away with scale, and obtaining competent embodied agents will instead require training directly on interaction data. Just as human proficiency in video games and physical tasks depends more on experience than intellectual capabilities, our position is that LLM agents will never be as capable as agents trained directly on interaction data, since they lack the common-sense behavioral priors necessary for low-level control.
In our proof-of-concept, we considered training an agent on general gameplay data before fine-tuning on a specific task. However, more generally, we could envision training a gaming foundation agent across a range of games (using our unified pixel observation space and gamepad controller action space), which could then be fine-tuned and aligned for a particular game. Similarly, while we primarily consider video games, we believe the modern LLM training paradigm (described more generally in Appendix A) could also be applied to training decision-making agents in other domains such as robotics. This could enable more general and robust agents, leveraging the impressive progress in LLMs while avoiding the fundamental limitations of LLM agents trained without real-world interaction data.
Finally, as LLMs inevitably continue to evolve, the developments can continue to be transferred to agents. For example, recent progress on improving LLM reasoning via process (lightman_lets_2023) or outcome rewards (lambert_tulu_2024; deepseek-ai_deepseek-r1_2025), or inference-time scaling (openai_openai_2024), can be incorporated into this framework simply as conventional RL with environment rewards and search. But by making this stage the ‘cherry-on-the-cake’ (lecun_nips_2016) rather than the sole training procedure, we can avoid many of the difficulties traditionally associated with RL, to obtain more general, robust and useful agents.
5 Conclusions
In this paper, we have taken the position that general agents should be trained like large language models. We have provided evidence for our position by following the LLM paradigm to train an agent from pixels to perform a desired behavior on a 3D AAA game as a proof-of-concept. This behavior would be difficult to achieve reliably with imitation learning or RL alone, demonstrating the value of each stage of the standard LLM training pipeline in the context of agents. We believe there is potential for many other recent developments in LLMs to transfer to decision-making agents, such as optimal dataset compositions, scaling laws, architectural and context improvements, post-training alignment procedures, and inference-time scaling. This will likely open up many exciting avenues of research in the coming years that have the potential to address many long-standing limitations and open-problems for decision-making agents. We hope our paper provides a framework for such research, and encourages communication and collaboration between the LLM and decision-making agent communities, to enable shared insights and provide a path towards more general and reliable agents for real-world applications.
Acknowledgements